Circular No. 8584

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005di IN MCG –04-52-46 Further to IAUC 8583, H. Pugh and W. Li report the LOSS discovery

of an apparent supernova on unfiltered KAIT images taken on Aug. 12.36 (at mag 18.6) and 13.36 UT (mag 18.3). The new object is located at α = 22^{h}25^{m}32^{s}.84, δ = −24^{o}14′20′′.1 (equinox 2000.0), which is 15′′.0 west and 9′′.2 north of the nucleus of MCG –04-52-46. A KAIT image taken on Aug. 3.32 showed nothing at this position (limiting mag ∼ 19.5).

M. Salvo, B. Schmidt, T. Davis, and S. Sankarankutty, Australian National University (ANU), report that a spectrogram (range 340–900 nm) of SN 2005di, obtained on Aug. 16.56 UT with the ANU 2.3-m telescope (+ Double-Beam Spectrograph), is similar to that of the type-Ia supernova 1994ae a few days past maximum light.

SUPERNOVA 2005df IN NGC 1559 Additional visual magnitude estimates by R. Evans (cf. IAUC 8580)

suggest that this object is again brightening: Aug. 9.8 UT, 13.0; 10.8, 12.8; 11.7, 13.0; 12.7, 13.0; 15.8, 12.8; 16.8, 12.5–12.6.

COMETS C/2005 N6, N7, N8, N9, N10, O3 (SOHO) Several more Kreutz sungrazing comets have been found on SOHO

images (cf. IAUC 8577). C/2005 N7–N10 appeared stellar in the C3 images (C/2005 N8 and C/2005 N10 being tiny, very faint, and tailless). C/2005 N7 reached mag ∼ 5.5 at 10.0R_{⊙} on July 12.612 UT in the C3 images, but was slightly diffuse and faded rapidly in the C2 images; C/2005 N9 reached mag ∼ 5.5 at 9.8R_{⊙} on July 14.012 in C3 images, but was slightly diffuse with an unmeasurable hint of tail in C2 images. C/2005 O3 showed a bright teardrop and an intermittent tail in C3 data, reaching mag ∼ 4.5 at 11.1R_{⊙} on July 23.138, while in C2 images it showed a very long, faint tail (mostly ‘headless’) that reached a length of 17′.2 at 5.8R_{⊙} on July 23.462.

Corrigendum. The reference to the observations and orbital elements for C/2005 N6 given on IAUC 8577 should read MPEC 2005-O35.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 N7 July 11.846 7^{h}08.^{m}8 +18^{o}56′ C3/2 RM 2005-Q03
C/2005 N8 12.638 7 13.3 +19 45 C3 BZ 2005-Q03
C/2005 N9  13.113 7 13.6 +18 26 C3/2 RM 2005-Q03
C/2005 N10 13.363 7 16.8 +19 16 C3 BZ 2005-Q03
C/2005 O3  21.971 7 43.2 +16 51 C3/2 RM 2005-Q03
2005 August 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green